a farmer has 800m of fencing material to enclose a rectangular field. the width of the field is 50m less than the length. find the dimensions of the field
Answer by Boreal(15235)

The width is x and the length is x+50.
the perimeter is 2width+2length
2x+2(x+50)=800
4x+100=800
4x=700
x=175 feet width
x+50=225 feet length